# GASP: Gradient-Aware Shortest Path Algorithm for Boundary-Confined 2-Manifold Reeb Graph Visualization

### v1

@@ 5 Implementation
In this study, we utilized triangular meshes as our input, $\mathcal{M}$}. For simplicity of the description, we show a principal direction as the height function, $f$, on the torus and modified torus models found in Fig.˜4-8. However, as we will describe, our algorithm is independent of this constraint, and our evaluation uses height and geodesic functions. To produce the Reeb graph, we utilized Recon [19].
@@ 5.1.1 Splitting Between Critical Points
The source and destination cut operations are then performed per triangle. For a source cut, triangles fall into one of 3 cases: (1) completely above the cut value, which is retained as is (see LABEL:fig:cut:src_before/LABEL:fig:cut:src_after type 1); (2) completely below the cut value, which is excluded from further computation (see LABEL:fig:cut:src_before/LABEL:fig:cut:src_after type 2); and (3) crossing the cut value, in which case triangles are cut along that value. The portion above the cut line is retained in the form of a smaller triangle (see LABEL:fig:cut:src_before/LABEL:fig:cut:src_after type 3a) or a quadrilateral, which is subdivided into 2 triangles (see LABEL:fig:cut:src_before/LABEL:fig:cut:src_after type 3b). A similar process is repeated for the destination cut, shown in LABEL:fig:cut:dst_before/LABEL:fig:cut:dst_after. LABEL:fig:decomp:src and LABEL:fig:decomp:dst show example source and destination cuts on models, respectively.
Figure 5: Examples of triangles (a/c) before and (b/d) after (a/b) source and (c/d) destination cuts. The 4 types of source cuts include: (1) completely above, (2) completely below; (3a) partially above keeping a triangle, and (3b) partially above keeping a quadrilateral. Beige triangles are retained, while gray ones are discarded.
@@ 5.3 Reeb Graph Assembly
For each pair of connected critical points, we generated the arcs of the Reeb graph as described above. Combining all of the arcs, the complete Reeb graph is realized. Fig.˜8 shows examples of the torus and modified torus with both the boundary and interior approach.
@@ 6 Evaluation
We have utilized 30 triangle meshes from [3]. All models were normalized such that their largest dimension range was [1,1][-1,1]. For each model, we generated Reeb graphs for height functions in all three principal directions (i.e., $x$, $y$, and $z$) and geodesic distance functions that emanate from the most extreme point in each principal direction (i.e., toptop, bottombottom, leftleft, rightright, backback, and frontfront). While the Reeb graph arc paths are calculated using our approach, all visualizations are rendered using ParaView [2]. Only a subset of the datasets are discussed in the paper, but the analysis of all datasets, source code, and a video of our approach compared to GB (TTK) can be found in the supplemental materials.
@@
As outlined in Section˜4.2.2, our approach features two main variations: boundary and interior. Both approaches incorporate a contour spacing parameter, $S$, while the interior approach has an additional buffer size parameter, $B$. The smaller the contour spacing, the finer the detail in the output. Throughout the evaluation, we utilize 3 levels of contour spacing with the largest being 0.1 (see LABEL:fig:duck-c-0.1), representing 5% of the [1,1][-1,1] domain, and 0.05 (see LABEL:fig:duck-c-0.05) and 0.025 (see LABEL:fig:duck-c-0.025) representing $2\times$ and $4\times$ the resolution, respectively. The buffer spacing determines the minimum distance an arc should be from the contour in the interior method. Again, 3 levels are utilized from 0 (see LABEL:fig:duck-i-0.00), which allows arcs to touch the contour, to 0.025 (see LABEL:fig:duck-i-0.025), to 0.05 (see LABEL:fig:duck-i-0.05) representing 2.5% of the domain. Parameters for testing were selected such that at the lowest level of detail, the output quality would be good but noticeably different from the highest level of detail. We selected the highest level of detail such that additional detail would not be noticeable.
@@ 6.1 Property 1: Arcs Confined to the Model
The first property for visualizing Reeb graphs is that arcs should ideally be contained within the model. For example, in LABEL:fig:i1-BvC:ttk:10-LABEL:fig:i1-BvC:ttk:15, the GB Reeb graph generated by TTK has several arcs outside of the horse mesh. The position of these arcs outside of the mesh is important, as it may misrepresent the function, which travels on the surface of the mesh, and cause unnecessary visual clutter.
@@
Figure 11: Examples for Property 1, arc confined to the model. Top: Examples comparing (a-b) GASP arcs that remain confined to the model’s boundary, while (c-d) some GB arcs go far beyond it. Bottom: (e-f) At lower contour spacing, the Reeb graph is more likely to remain inside the model, (g) while at larger contour spacing values, arcs are more likely to extend outside the model (e.g., near the neck). (h) When buffer size is 0, arcs around the front legs are partially outside the boundary (not easily visible in the figures). (i-j) When buffer size increases to $0.025$ and $0.05$, arcs move further inside the model. (k) One GASP limitation is that for critical points with small function distance and large Euclidean distance, long arcs can appear. (l) Interestingly and somewhat counterintuitively, as contour spacing increases (i.e., fewer contours are used), these long arcs disappear because topological cylinders switch from regular features to thin features. Complete Reeb graphs include red and blue arcs; red arcs highlight comparisons. Additional examples in supplement.
@@
GASP uses a strategies that try to constrain the Reeb graph to the mesh surface using the boundary approach (see LABEL:fig:i1-BvC:ours:boundary) or inside the mesh when using the interior approach (see LABEL:fig:i1-BvC:ours:interior), while GB Reeb graphs extend beyond the boundary (see LABEL:fig:i1-BvC:ttk:10-LABEL:fig:i1-BvC:ttk:15). We quantitatively evaluate the effectiveness of GASP and GB by measuring the ratio of arc $Length_{outside}/Length_{Total}$ (0 means the arc is entirely interior, while $1$ indicates the arc is entirely outside the mesh). The ratio is calculated per arc, and we present the average values per Reeb graph. To measure how far an arc extends beyond the model, we calculate the sum of the areas of all arc segments outside the mesh surface. In particular, if an arc segment of length $l$ whose endpoints are $d_{1}$ and $d_{2}$ units away from the boundary of the mesh, respectively, the area is $a={l}\cdot\frac{d_{1}+d_{2}}{2}$. Lower area values indicate the extended arcs are closer to the mesh. The scatter plots in LABEL:fig:vis-i1-CwT-x (ratio of arcs outside the model) and LABEL:fig:vis-i1-area-CwT-x (area of arcs outside the model) show the improved performance of GASP compared to the two TTK variants of GB across different models for the same function (scatter plots for additional functions are provided in the supplement). Furthermore, the GASP interior method is generally more effective than the boundary approach.
@@ 6.2 Property 2: Compact Arcs
When TTK creates a Reeb graph of the model using GB, it transforms the straight lines into arcs by utilizing arc sampling and smoothing. The length of the arcs increases as the arc sampling parameter increases. Moreover, in some cases, adjusting the iteration parameter allows the arc to position itself within the model at the cost of increased length. Fig.˜10 shows an example where increasing the sampling causes the arcs to respect the model boundary while also getting longer. In contrast, GASP employs the shortest path to ensure that the path taken is minimal while also respecting the boundary (up to the limits of contour spacing). An example is illustrated in LABEL:fig:i2:BvC:ours:boundary-LABEL:fig:i2:BvC:ttk:15. GB results in several unnecessarily long arcs (highlighted in red), with the less smooth version (see LABEL:fig:i2:BvC:ttk:10) producing shorter arcs than the smoother version (see LABEL:fig:i2:BvC:ttk:15). However, both variations of GASP, respect the boundary of the model while keeping their length minimal (see LABEL:fig:i2:BvC:ours:boundary-LABEL:fig:i2:BvC:ours:interior). To quantitatively measure how well a Reeb graph minimizes arc length, we measure the total length of each Reeb graph arc and compare it to the Euclidean distance between the associated critical points. The ratio $Length_{Total}/Distance_{CP}$ are calculated per arc, and smaller is generally better ($1$ is a straight edge). We report the average values per Reeb graph. LABEL:fig:issue2:BvC shows that for almost all cases, GASP arcs are significantly shorter than GB arcs and that the GASP interior approach arcs are shorter than boundary approach arcs.
@@
Our evaluation shows that GASP generally produces shorter arcs than GB, and increases in both contour spacing and buffer size generally result in small increases in arc length. One important note about these results is that GASP does particularly well on nearby critical points because it connects the critical points more directly. However, for some longer arcs, particularly in non-convex areas of high curvature, pose challenges to GASP that may result in producing longer arcs because the model’s boundary constrains it. LABEL:fig:i2:Limitation:TTK shows an example where the GB arcs in red are shorter than the GASP arcs in LABEL:fig:i2:Limitation:ours-boundary.
@@ 6.3 Property 3: Following the Function Gradient
GASP creates arcs that move contour to contour, meaning that its Reeb graphs move with the gradient, such as in LABEL:fig:i3:BvC:ours:boundary-LABEL:fig:i3:BvC:ours:interior. On the other hand, the sampling and smoothing of GB (TTK) cause its Reeb graph arcs (highlighted in red) to move up and down (LABEL:fig:i3:BvC:ttk:10-LABEL:fig:i3:BvC:ttk:15), nearly orthogonal to the function gradient (from left to right). To measure how well the direction of Reeb graph arcs reflect the flow direction of the function, we measure the total length in the principal direction of the function (i.e., $x$, $y$, or $z$) and compare it to the total length in non-principal directions (i.e., $y+z$, $x+z$, and $x+y$, respectively) [^2] For ease of calculating the principal and non-principal direction, we only perform this analysis on height functions.. Values are calculated per arc using the ratio of $(Distance_{NP_{1}}+Distance_{NP_{2}})/Length_{PD}$, where $PD$ is the principal direction and $NP_{1}$ and $NP_{2}$ are the non-principal directions (lower is better). The results in LABEL:fig:issue3:BvC show that for most models, GASP produced arcs better aligned with the gradient than GB.
@@
Figure 16: Quantitative analysis for Property 3, following function gradient, using mean ratio of arc length in non-principal directions to the principal direction (log scale, lower is better). (a) Comparison of GASP and GB across all models shows that for the majority of configurations, GASP approaches yield better alignment with the gradient compared to both GB variations. (b) Charts illustrating the effect of contour spacing on GASP for the subset of models used in the paper. The near 0 slope indicates a negligible impact on the alignment arcs with the gradient. (c) Charts illustrating the effect of buffer size on GASP for the same subset of models. The near 0 slope indicates a negligible impact on the alignment arcs with the gradient. Additional examples in supplement.
@@ 6.4 Property 4: Optional Arc Smoothness
The arc smoothness of the Reeb graph is an aesthetic quality that can provide the auxiliary benefit of a clearer representation [37, 15]. However, ideally, improvements in smoothness should be achieved without sacrificing the three faithfulness properties. Our primary goal was to improve the faithfulness of the Reeb graph visualization, a choice which necessitates a compromise in this aesthetic aspect. This trade-off is observed as lower smoothness in some GASP Reeb graphs. To quantify this, we calculated smoothness, $M$, as the mean of the tangents of a Reeb graph arc, with a corresponding value $\alpha=\frac{1}{1+M}$ in the range $[0,1]$ (lower is smoother). LABEL:fig:Smoothness compares the smoothness of GASP and GB for the models. The GASP boundary approach achieves better smoothness than the interior approach, as interior arcs must deviate from the boundary and critical points, leading to more abrupt path changes (see LABEL:fig:smoothness:ours:boundary-regular-LABEL:fig:smoothness:ours:interior-regular), whereas GB produces smoother Reeb graphs for nearly all models (see LABEL:fig:smoothness:ttk5-LABEL:fig:smoothness:ttk15).
@@ 6.5 Computational Performance
To experimentally confirm this analysis, we recorded the execution time of all models and function directions on a M1 Mac Mini for the boundary approach (see LABEL:fig:time:boundary) and interior approach (see LABEL:fig:time:interior) across different contour spacing and buffer size. Reeb graph arcs were calculated using a thread pool of 8 simultaneous threads. The resulting performance was linear with respect to the total number of triangles in the mesh times the number of Reeb graph edges (goodness of fit, $R^{2}\approx 0.8$ for all boundary configurations, and $R^{2}\approx 0.7$ for all interior configurations). The performance was slightly lower as contour spacing was reduced (i.e., more contours) and as buffer size for the interior approach was reduced (i.e., more points inside the contours).
@@
Comparing the computational performance of GASP to TTK implementation of GB is an apple-to-oranges comparison, as GASP is implemented as a standalone in Python, and TTK utilizes C++ implementation of GB as part of Paraview. Nevertheless, for context, LABEL:fig:time:ttk compares the computational performance of GB with regression lines for several variations of GASP’s boundary and interior approaches. GB’s performance was, unsurprisingly, better than GASP. We also observed that GB’s performance was largely parameter dependent, not data dependent.

### v2

@@ 5 Implementation
In this study, we utilized triangular meshes as our input, $\mathcal{M}$}. To simplify the description, we show a principal direction as the height function, $f$, on the torus and modified torus models found in Fig.˜4-8. However, as we will describe, our algorithm is independent of this constraint, and our evaluation uses height and geodesic functions. To produce the Reeb graph, we utilized Recon [19].
@@ 5.1.1 Splitting Between Critical Points
The source and destination cut operations are then performed per triangle. For a source cut, triangles fall into one of three cases: (1) completely above the cut value, which is retained as is (see LABEL:fig:cut:src_before/LABEL:fig:cut:src_after type 1); (2) completely below the cut value, which is excluded from further computation (see LABEL:fig:cut:src_before/LABEL:fig:cut:src_after type 2); and (3) crossing the cut value, in which case triangles are cut along that value. The portion above the cut line is retained in the form of a smaller triangle (see LABEL:fig:cut:src_before/LABEL:fig:cut:src_after type 3a) or a quadrilateral, which is subdivided into two triangles (see LABEL:fig:cut:src_before/LABEL:fig:cut:src_after type 3b). A similar process is repeated for the destination cut, shown in LABEL:fig:cut:dst_before/LABEL:fig:cut:dst_after. LABEL:fig:decomp:src and LABEL:fig:decomp:dst show example source and destination cuts on models, respectively.
Figure 5: Examples of triangles (a/c) before and (b/d) after (a/b) source and (c/d) destination cuts. The four types of source cuts include: (1) completely above, (2) completely below, (3a) partially above keeping a triangle, and (3b) partially above keeping a quadrilateral. Beige triangles are retained, while gray ones are discarded.
@@ 5.3 Reeb Graph Assembly
For each pair of connected critical points, we generated the arcs of the Reeb graph as described above. Combining all the arcs, the complete Reeb graph is realized. Fig.˜8 shows examples of the torus and modified torus with both the boundary and interior approach.
@@ 6 Evaluation
We have utilized 30 triangle meshes from [3]. All models were normalized such that their largest dimension range was [1,1][-1,1]. For each model, we generated Reeb graphs for height functions in all three principal directions (i.e., $x$, $y$, and $z$) and geodesic distance functions that emanate from the most extreme point in each principal direction (i.e., toptop, bottombottom, leftleft, rightright, backback, and frontfront). While the Reeb graph arc paths are calculated using our approach, all visualizations are rendered using ParaView [2]. Only a subset of the datasets is discussed in the paper, but the analysis of all datasets, source code, and a video of our approach compared to GB (TTK) can be found in the supplemental materials.
@@
As outlined in Section˜4.2.2, our approach features two main variations: boundary and interior. Both approaches incorporate a contour spacing parameter, $S$, while the interior approach has an additional buffer size parameter, $B$. The smaller the contour spacing, the finer the detail in the output. Throughout the evaluation, we utilize three levels of contour spacing with the largest being 0.1 (see LABEL:fig:duck-c-0.1), representing 5% of the [1,1][-1,1] domain, and 0.05 (see LABEL:fig:duck-c-0.05) and 0.025 (see LABEL:fig:duck-c-0.025) representing $2\times$ and $4\times$ the resolution, respectively. The buffer spacing determines the minimum distance an arc should be from the contour in the interior method. Again, three levels are utilized from 0 (see LABEL:fig:duck-i-0.00), which allows arcs to touch the contour, to 0.025 (see LABEL:fig:duck-i-0.025), to 0.05 (see LABEL:fig:duck-i-0.05) representing 2.5% of the domain. Parameters for testing were selected such that at the lowest level of detail, the output quality would be good but noticeably different from the highest level of detail. We selected the highest level of detail such that additional detail would not be noticeable.
@@ 6.1 Property 1: Arcs Confined to the Model
The first property for visualizing Reeb graphs is that arcs should ideally be contained within the model. For example, in LABEL:fig:i1-BvC:ttk:10-LABEL:fig:i1-BvC:ttk:15, the GB Reeb graph generated by TTK has several arcs outside of the horse mesh. The position of these arcs outside of the mesh is important, as it may misrepresent the function, which travels on the surface of the mesh, and causes unnecessary visual clutter.
@@
Figure 11: Examples for Property 1, arcs confined to the model. Top: Examples comparing (a-b) GASP arcs that remain confined to the model’s boundary, while (c-d) some GB arcs go far beyond it. Bottom: (e-f) At lower contour spacing, the Reeb graph is more likely to remain inside the model, (g) while at larger contour spacing values, arcs are more likely to extend outside the model (e.g., near the neck). (h) When buffer size is 0, arcs around the front legs are partially outside the boundary (not easily visible in the figures). (i-j) When buffer size increases to $0.025$ and $0.05$, arcs move further inside the model. (k) One GASP limitation is that for critical points with small function distance and large Euclidean distance, long arcs can appear. (l) Interestingly and somewhat counterintuitively, as contour spacing increases (i.e., fewer contours are used), these long arcs disappear because topological cylinders switch from regular features to thin features. Complete Reeb graphs include red and blue arcs; red arcs highlight comparisons. Additional examples in supplement.
@@
GASP uses strategies that try to constrain the Reeb graph to the mesh surface using the boundary approach (see LABEL:fig:i1-BvC:ours:boundary) or inside the mesh when using the interior approach (see LABEL:fig:i1-BvC:ours:interior), while GB Reeb graphs extend beyond the boundary (see LABEL:fig:i1-BvC:ttk:10-LABEL:fig:i1-BvC:ttk:15). We quantitatively evaluate the effectiveness of GASP and GB by measuring the ratio of arc $Length_{outside}/Length_{Total}$ (0 means the arc is entirely interior, while $1$ indicates the arc is entirely outside the mesh). The ratio is calculated per arc, and we present the average values per Reeb graph. To measure how far an arc extends beyond the model, we calculate the sum of the areas of all arc segments outside the mesh surface. In particular, if an arc segment of length $l$ whose endpoints are $d_{1}$ and $d_{2}$ units away from the boundary of the mesh, respectively, the area is $a={l}\cdot\frac{d_{1}+d_{2}}{2}$. Lower area values indicate the extended arcs are closer to the mesh. The scatter plots in LABEL:fig:vis-i1-CwT-x (ratio of arcs outside the model) and LABEL:fig:vis-i1-area-CwT-x (area of arcs outside the model) show the improved performance of GASP compared to the two TTK variants of GB across different models for the same function (scatter plots for additional functions are provided in the supplement). Furthermore, the GASP interior method is generally more effective than the boundary approach.
Mean Arc Outside Ratio
Arc Area Outside
@@ 6.2 Property 2: Compact Arcs
When TTK creates a Reeb graph of the model using GB, it transforms the straight lines into arcs by utilizing arc sampling and smoothing. The length of the arcs increases as the arc sampling parameter increases. Moreover, in some cases, adjusting the iteration parameter allows the arc to position itself within the model at the cost of increased length. Fig.˜10 shows an example where increasing the sampling causes the arcs to respect the model boundary while also getting longer. In contrast, GASP employs the shortest path to ensure that the path taken is minimal while also respecting the boundary (up to the limits of contour spacing). An example is illustrated in LABEL:fig:i2:BvC:ours:boundary-LABEL:fig:i2:BvC:ttk:15. GB results in several unnecessarily long arcs (highlighted in red), with the less smooth version (see LABEL:fig:i2:BvC:ttk:10) producing shorter arcs than the smoother version (see LABEL:fig:i2:BvC:ttk:15). However, both variations of GASP, respect the boundary of the model while keeping their length minimal (see LABEL:fig:i2:BvC:ours:boundary-LABEL:fig:i2:BvC:ours:interior). To quantitatively measure how well a Reeb graph minimizes arc length, we measure the total length of each Reeb graph arc and compare it to the Euclidean distance between the associated critical points. The ratio $Length_{Total}/Distance_{CP}$ is calculated per arc, and smaller is generally better ($1$ is a straight edge). We report the average values per Reeb graph. LABEL:fig:issue2:BvC shows that for almost all cases, GASP arcs are significantly shorter than GB arcs and that the GASP interior approach arcs are shorter than boundary approach arcs.
@@
Our evaluation shows that GASP generally produces shorter arcs than GB, and increases in both contour spacing and buffer size generally result in small increases in arc length. One important note about these results is that GASP does particularly well on nearby critical points because it connects the critical points more directly. However, some longer arcs, particularly in non-convex areas of high curvature, pose challenges to GASP that may result in producing longer arcs because the model’s boundary constrains it. LABEL:fig:i2:Limitation:TTK shows an example where the GB arcs in red are shorter than the GASP arcs in LABEL:fig:i2:Limitation:ours-boundary.
@@ 6.3 Property 3: Following the Function Gradient
GASP creates arcs that move contour to contour, meaning that its Reeb graphs move with the gradient, such as in LABEL:fig:i3:BvC:ours:boundary-LABEL:fig:i3:BvC:ours:interior. On the other hand, the sampling and smoothing of GB (TTK) cause its Reeb graph arcs (highlighted in red) to move up and down (LABEL:fig:i3:BvC:ttk:10-LABEL:fig:i3:BvC:ttk:15), nearly orthogonal to the function gradient (from left to right). To measure how well the direction of Reeb graph arcs reflects the flow direction of the function, we measure the total length in the principal direction of the function (i.e., $x$, $y$, or $z$) and compare it to the total length in non-principal directions (i.e., $y+z$, $x+z$, and $x+y$, respectively) [^2] For ease of calculating the principal and non-principal direction, we perform this analysis only on height functions.. Values are calculated per arc using the ratio of $(Distance_{NP_{1}}+Distance_{NP_{2}})/Length_{PD}$, where $PD$ is the principal direction and $NP_{1}$ and $NP_{2}$ are the non-principal directions (lower is better). The results in LABEL:fig:issue3:BvC show that for most models, GASP produced arcs better aligned with the gradient than GB.
@@
Figure 16: Quantitative analysis for Property 3, following function gradient, using the mean ratio of arc length in non-principal directions to the principal direction (log scale, lower is better). (a) Comparison of GASP and GB across all models shows that for the majority of configurations, GASP approaches yield better alignment with the gradient compared to both GB variations. (b) Charts illustrating the effect of contour spacing on GASP for the subset of models used in the paper. The near 0 slope indicates a negligible impact on the alignment arcs with the gradient. (c) Charts illustrating the effect of buffer size on GASP for the same subset of models. The near 0 slope indicates a negligible impact on the alignment arcs with the gradient. Additional examples in the supplement.
@@ 6.4 Property 4: Optional Arc Smoothness
The arc smoothness of the Reeb graph is an aesthetic quality that can provide the auxiliary benefit of a clearer representation [37, 15]. However, ideally, improvements in smoothness should be achieved without sacrificing the three faithfulness properties. Our primary goal was to improve the faithfulness of the Reeb graph visualization, a choice that necessitates a compromise in this aesthetic aspect. This trade-off is observed as lower smoothness in some GASP Reeb graphs. To quantify this, we calculated smoothness, $M$, as the mean of the tangents of a Reeb graph arc, with a corresponding value $\alpha=\frac{1}{1+M}$ in the range $[0,1]$ (lower is smoother). LABEL:fig:Smoothness compares the smoothness of GASP and GB for the models. The GASP boundary approach achieves better smoothness than the interior approach, as interior arcs must deviate from the boundary and critical points, leading to more abrupt path changes (see LABEL:fig:smoothness:ours:boundary-regular-LABEL:fig:smoothness:ours:interior-regular), whereas GB produces smoother Reeb graphs for nearly all models (see LABEL:fig:smoothness:ttk5-LABEL:fig:smoothness:ttk15).
@@ 6.5 Computational Performance
To experimentally confirm this analysis, we recorded the execution time of all models and function directions on a M1 Mac Mini for the boundary approach (see LABEL:fig:time:boundary) and interior approach (see LABEL:fig:time:interior) across different contour spacing and buffer size. Reeb graph arcs were calculated using a thread pool of eight simultaneous threads. The resulting performance was linear with respect to the total number of triangles in the mesh times the number of Reeb graph edges (goodness of fit, $R^{2}\approx 0.8$ for all boundary configurations, and $R^{2}\approx 0.7$ for all interior configurations). The performance was slightly lower as contour spacing was reduced (i.e., more contours) and as the buffer size for the interior approach was reduced (i.e., more points inside the contours).
@@
Comparing the computational performance of GASP to the TTK implementation of GB is an apples-to-oranges comparison, as GASP is implemented as a standalone in Python, and TTK utilizes C++ implementation of GB as part of Paraview. Nevertheless, for context, LABEL:fig:time:ttk compares the computational performance of GB with regression lines for several variations of GASP’s boundary and interior approaches. GB’s performance was, unsurprisingly, better than GASP. We also observed that GB’s performance was largely parameter dependent, not data dependent.